OpenStudy (anonymous):

The sum of two consecutive integers is 141. Find the integers. a.n+1=141 b.2n+1=141 c.2n+2=141 d.2n=141 (

OpenStudy (anonymous):

(a number)+(a number +1) = 141

OpenStudy (anonymous):

b

OpenStudy (anonymous):

did anyone choose c.

OpenStudy (anonymous):

n = 70 (2 x 70 ) + 1 =141